# RealCamo: Boosting Real Camouflage Synthesis with Layout Controls and Textual-Visual Guidance

### v1

@@ Abstract
Camouflaged image generation (CIG) has recently emerged as an efficient alternative for acquiring high-quality training data for camouflaged object detection (COD). However, existing CIG methods still suffer from a substantial gap to real camouflaged imagery: generated images either lack sufficient camouflage due to weak visual similarity, or exhibit cluttered backgrounds that are semantically inconsistent with foreground targets. To address these limitations, we propose ReamCamo, a unified out-painting based framework for realistic camouflaged image generation. ReamCamo explicitly introduces additional layout controls to regulate global image structure, thereby improving semantic coherence between foreground objects and generated backgrounds. Moreover, we construct a multi-modal textual–visual condition by combining a unified fine-grained textual task description with texture-oriented background retrieval, which jointly guides the generation process to enhance visual fidelity and realism. To quantitatively assess camouflage quality, we further introduce a background–foreground distribution divergence metric that measures the effectiveness of camouflage in generated images. Extensive experiments and visualizations demonstrate the effectiveness of our proposed framework.
Code and results will be released soon.
Camouflaged Image Generation, Diffusion Model, Controlled Image Generation, Synthesize Data
Figure 1: Motivation and qualitative comparison of camouflaged image synthesis. (a) LAKE-RED (CVPR2024LAKERED) fails to achieve effective camouflage, with targets remaining visually distinguishable, whereas our method produces images with well-camouflaged targets. (b) CamoAny (CVPR2025CamoAny) achieves visual camouflage but suffers from weak semantic consistency between the target and the generated background, in contrast, our method preserves semantic coherence, resulting in more realistic scenes. (c) Our approach is capable of generating diverse and realistic camouflaged images while jointly satisfying visual similarity and semantic consistency.
@@ 1 Introduction
Why Synthesize Data: Concealed visual perception tasks, exemplified by camouflaged object detection (COD) (ICML2025RUN), aim to identify semantically meaningful targets that are visually indistinguishable from their surrounding backgrounds. Such tasks are increasingly important in real-world applications, including medical diagnosis, agricultural and industrial production (AIR2024CODsurvey). Despite recent progress, the development of COD remains fundamentally constrained by the limited scale of available training data.
@@
In contrast to large-scale benchmarks for general object detection & segmentation, such as PASCAL VOC (IJCV2010PASCAL) and MS-COCO (ECCV2014MScoco), which contain tens to hundreds of thousands of precisely annotated images, existing COD datasets (CVPR2020SINet-COD10K; CVPR2021LSR-NC4K) typically comprise only a few thousand samples. This scarcity arises from two intrinsic challenges: (i) camouflaged objects are inherently difficult to annotate due to its high visual similarity to backgrounds, and (ii) camouflaged scenes with high quality are themselves rare and costly to capture in the real world. As a result, data insufficiency has become a critical bottleneck for improving both the robustness and generalization of COD models.
@@
Recent advances in image generation models, including GAN (ACM2020GAN) and diffusion models (NIPS2020DDPM; CVPR2022LDM), offer a promising alternative, while leveraging high-quality synthetic data to augment training has been shown to be an effective and cost-efficient strategy in multiple vision tasks such as salient object detection (arXiv2025S3OD). This observation motivates the camouflaged image synthesis task, whose core objective is to generate realistic and diverse camouflaged scenes that can meaningfully complement limited real-world datasets and thereby improve downstream COD performance.
@@
Figure 2: The proposed KL divergence based metric. Existing generative evaluation metrics, such as FID and KID, are insufficient to measure the similarity between foreground and background, while our proposed $\mathrm{KL}_{BF}$ provides a quantitative assessment of camouflage. Top: Image with a higher $\mathrm{KL}_{BF}$. Bottom: Our generated image with a lower $\mathrm{KL}_{BF}$, more camouflaged.
@@
Why Out-Painting: Existing approaches to camouflaged image synthesis can be broadly categorized into two paradigms. The first paradigm, exemplified by style-transfer based methods such as LCGNet (TMM2022LCGNet), modifies the appearance of a target object to match a given background. However, this process often alters object-specific textures and visual cues to such an extent that the semantic identity of the target is partially or entirely compromised. Consequently, these methods achieve camouflage by enforcing visual uniformity, rather than by modeling the natural interaction between an object and its environment.
@@
The second paradigm, represented by out-painting based methods such as LAKE-RED (CVPR2024LAKERED), follows a fundamentally different strategy. Instead of altering the target object, these methods preserve its semantic content and generate surrounding background regions whose textures and structures are consistent with the target appearance. This formulation enables camouflage to emerge from contextual compatibility between the object and its environment, rather than from direct manipulation of the object itself.
@@
Compared with style-transfer based approaches, out-painting offers several advantages. First, it maintains the semantic integrity of the target while avoiding the need for an externally provided background image. More importantly, it aligns more closely with the natural principles of camouflage observed in the real world. Camouflage is inherently a relative and environment-dependent phenomenon: the same object may appear well camouflaged in one context but highly salient in another (arXiv2025COD-SOD). From the perspective of natural selection, creatures do not actively redesign their appearance to match each environment, instead, environmental variation favors the survival of individuals whose appearances are already compatible with their surroundings. This perspective naturally corresponds to the out-painting formulation.
@@
Why Background Control: For out-painting based methods to produce realistic and effective camouflaged training data, we argue that two key requirements must be simultaneously satisfied. First, the generated background should exhibit strong appearance compatibility with the target, particularly in terms of color distribution and texture patterns, so that the target is genuinely camouflaged under visual perception. Second, the background must convey semantically plausible content consistent with real-world priors, for instance, aquatic objects should appear in water rather than in visually incompatible contexts.
@@
(1) Controlled camouflaged image synthesis. To overcome the limitations of existing out-painting based methods in background generation, we introduce a unified out-painting based framework termed RealCamo, which integrates explicit layout controls and textual-visual guidance. We argue that background layout plays a critical role in determining semantic plausibility. Building upon the contrast control in CamoAny, we incorporate additional depth and holistically-nested edge information into ControlNet to explicitly constrain the spatial layout of generated scenes, thereby preserving semantic realism. Furthermore, to enhance camouflage effectiveness at the texture level, we propose a textual–visual guidance mechanism that combines a unified fine-grained textual task descriptions with additional visual cues retrieved according to target texture characteristics, enhancing visual fidelity and realism of generation.
(2) Camouflage effectiveness evaluation metrics. A realistic camouflaged image should preserve the original scene layout while achieving camouflage through modifications in background color and texture. Existing generative evaluation metrics, such as FID and KID, are insufficient to capture this requirement. We therefore introduce SSIM to quantify structural consistency between generated and original images as a indicator for semantic realism. In addition, we propose a KL divergence based metric $\mathrm{KL}_{BF}$, which measures the similarity between background and foreground pixel distributions, providing a principled quantitative assessment of camouflage effectiveness, as shown in Fig. 2.
(3) Downstream evaluation. To assess the practical utility of the generated camouflaged images, we construct a large-scale synthetic dataset, SynCOD12K, as supplementary training data for camouflaged object detection. Extensive experiments on four COD benchmarks demonstrate consistent performance improvements, with further investigations on the effects of camouflage-augmented data on related salient and general object detection tasks.
@@ 2 Related Work
Synthetic Dataset Generation. Synthetic data has become an effective strategy for alleviating data scarcity in computer vision (MIR2024Synsurvey). Advances in generative models, including GANs and diffusion models, have enabled the construction of large-scale, high-quality synthetic datasets. DatasetGAN (CVPR2021DatasetGAN) generates image-annotation pairs by exploiting the latent feature space of pretrained GANs, while BigDatasetGAN (CVPR2022BigDatasetGAN) extends this paradigm to ImageNet-scale class diversity using VQGAN (CVPR2021VQGAN). More recently, diffusion-based methods such as DiffuMask (ICCV2023DiffuMask) and DatasetDM (CVPR2022LDM) leverage text-to-image models to produce photorealistic images and corresponding annotations for downstream perception tasks. Despite their success, these approaches primarily focus on generic image synthesis and annotation, and remain limited when applied to structurally complex and semantically subtle scenarios such as camouflaged image generation.
@@
Camouflaged Image Generation. Camouflaged image generation, introduced by TOG2010Chu, can be categorized into two paradigms: Style transfer based background-guided approaches modify foreground appearance to blend targets into given backgrounds. DCI (AAAI2020DCI) employs an attention-aware camouflage loss to suppress salient cues while preserving subtle perceptual features, and LCGNet (TMM2022LCGNet) achieves extreme camouflage by fusing high-level foreground and background representations. In contrast, out-painting based foreground-guided methods generate backgrounds conditioned on the foreground target. LAKE-RED (CVPR2024LAKERED) introduces a knowledge retrieval–augmented framework using a pretrained codebook, while FACIG (ICME2025FACIG) refines feature integration to reduce foreground distortion CamoAny (CVPR2025CamoAny) further improves background synthesis through controlled out-painting and representation engineering. Despite these advances, existing methods still struggle to simultaneously achieve strong visual similarity and semantic consistency. More recently, the text-driven method CT-CIG (AAAI2026CTCIG) has shown promise in realistic camouflage synthesis, but the inability to provide aligned annotations limits their applicability to downstream camouflaged object detection applications.
@@
Controlled Image Generation. Recent advances in text-to-image (T2I) generation integrate large-scale text encoders, such as CLIP (ICML2021CLIP) and T5 (JMLR2020T5), into diffusion models, yielding powerful T2I models. To enable fine-grained spatial control, ControlNet (ICCV2023ControlNet) introduces an auxiliary trainable branch to inject conditional signals into pretrained diffusion models. Building upon this design, UniControl (NIPS2023UniControl) employs a task-aware hypernetwork to modulate zero convolutions of the ControlNet in handling multiple conditions, while CtrLoRA (ICLR2025CtrLoRA) further reduces parameter overhead through condition-specific LoRA (ICLR2022LoRA) fine-tuning, and ICCV2025PreserveAnything proposes an N-channel ControlNet to improve object preservation, prompt alignment, and aesthetic quality under multiple control settings.
@@ 3.1 Preliminary
Our proposed method is based on the classic Latent Diffusion Models (LDMs) (CVPR2022LDM) and ControlNet (ICCV2023ControlNet). Here, we briefly revisiting their basics, with more details can be found in Appendix B.
Figure 3: The pipeline of our camouflaged image generation framework RealCamo. Explicit layout controls (i.e., the contrast, depth, and HED controls) are extracted by the Layout Controls Generator (LCG) and injected to ControlNet, while the multi-modal condition produced by the Textual-Visual Condition Generator (TVCG) is used as a guidance in refining the camouflage generation process.
Latent Diffusion Models. LDMs generate high-quality images by reversing a Markov forward process in the latent space. Specifically, Given a VAE (ICML2014VAE2) with encoder $\mathcal{E}$ and decoder $\mathcal{D}$, Gaussian noise $\epsilon$ is progressively added to a latent representation $z_{0}=\mathcal{E}(x_{0})$ of the original image $x_{0}$, resulting in a noisy image $z_{t}=\alpha_{t}z_{0}+(1-\alpha_{t})\epsilon$, with more noise being introduced when $\alpha_{t}$ decreases as timestep $t$ increases. A denoising network $\epsilon_{\theta}$ is trained to predict the added noise given a condition $c^{t}$ following:
@@
ControlNet. ControlNet extends the structural controllability of LDMs by incorporating addition conditions ($c^{f}$) like Canny edges and human pose. With the introduction of additional trained branch aided by zero convolutions, ControlNet reformulates Equation 1 as:
@@ 3.2 Framework Overview
As illustrated in Fig. 3, the proposed RealCamo framework is built upon Stable Diffusion 1.5 (CVPR2022LDM) and ControlNet (ICCV2023ControlNet), enabling controllable camouflaged image generation through complementary structural and semantic conditioning. Control signals are injected in two ways: (i) structural constraints introduced via ControlNet to regulate spatial layout, and (ii) multi-modal prompts injected through cross-attention to guide image content. These mechanisms align with our objective that high-quality camouflage requires both visual similarity and semantic consistency between foreground targets and background scenes. Accordingly, we employ a Layout Controls Generator to provide explicit structural priors via ControlNet, and a Textual–Visual Condition Generator that integrates a unified fine-grained task description with background images retrieved by a Texture-Oriented Background Retrieval module, producing informative multi-modal prompts to effectively steer the camouflage generation process.
@@ 3.3 Layout Controls Generator
We argue that semantic consistency in camouflaged image generation is largely determined by the structural layout of the image. Prior methods (CVPR2024LAKERED; CVPR2025CamoAny) often yield cluttered or physically implausible backgrounds due to the lack of explicit layout constraints. To address this, we propose a layout conditions generator (LCG) that extracts complementary structural cues and injects them via ControlNet, as shown in Fig. 3. Specifically, for a given image $\mathrm{I}$ and a corresponding mask $\mathrm{M}$ which indicate the foreground target, we employ LCG to extract contrast, depth, and holistically-nested edge (HED) controls as explicit structural priors to promote coherent spatial organization and improved foreground–background semantic alignment:
$c_{cst},c_{dep},c_{hed}=LCG(\mathrm{I},\mathrm{M},is\_training).$ (3)
Each control signal is embedded using a dedicated MLP (NIPS2023UniControl) due to distributional differences, and then fused as a structural guidance for ControlNet to process.
Contrast Control. Following CVPR2025CamoAny, contrast control $c_{cst}$ aims to preserve the foreground target while providing basic regulation of color contrast, varies when training and inference based on is_trainingis\_training.
Depth Control. Depth control $c_{dep}$ is generated using Depth Anything V2 (NIPS2024DepthAnythingV2), and is intended to constrain the spatial relationships within the image.
HED Control. HED control $c_{hed}$ is obtained by deploying Holistically-nested edge detection (ICCV2015HED) to regulate the contours of the main elements in the image.

### v2

@@ Abstract
Camouflaged image generation (CIG) has recently emerged as an efficient alternative for acquiring high-quality training data for camouflaged object detection (COD). However, existing CIG methods still suffer from a substantial gap to real camouflaged imagery: generated images either lack sufficient camouflage due to weak visual similarity, or exhibit cluttered backgrounds that are semantically inconsistent with foreground targets. To address these limitations, we propose RealCamo, a novel out-painting-based framework for controllable realistic camouflaged image generation. RealCamo explicitly introduces additional layout controls to regulate global image structure, thereby improving semantic coherence between foreground objects and generated backgrounds. Moreover, we construct a multimodal textual-visual condition by combining a unified fine-grained textual task description with texture-oriented background retrieval, which jointly guides the generation process to enhance visual fidelity and realism. To quantitatively assess camouflage quality, we further introduce a background-foreground distribution divergence metric that measures the effectiveness of camouflage in generated images. Extensive experiments and visualizations demonstrate the effectiveness of our proposed framework.
Code and results will be released.
Camouflaged Image Generation, Diffusion Model, Controlled Image Generation, Synthetic Data
Figure 1: Motivation and qualitative comparison of camouflaged image synthesis. (a) LAKE-RED (Zhao et al., 2024) fails to achieve effective camouflage, with targets remaining visually distinguishable, whereas our method produces images with well-camouflaged targets. (b) CamoAny (Das and Gopalakrishnan, 2025) achieves visual camouflage but suffers from weak semantic consistency between the target and the generated background; in contrast, our method preserves semantic coherence, resulting in more realistic scenes. (c) Our approach is capable of generating diverse and realistic camouflaged images while jointly satisfying visual similarity and semantic consistency.
@@ 1 Introduction
Why Synthetic Data: Concealed visual perception tasks, exemplified by camouflaged object detection (COD) (He et al., 2025), aim to identify semantically meaningful targets that are visually indistinguishable from their surrounding backgrounds. Such tasks are increasingly important in real-world applications, including medical diagnosis, agricultural and industrial production (Xiao et al., 2024). Despite recent progress, the development of COD remains fundamentally constrained by the limited scale of available training data.
@@
In contrast to large-scale benchmarks for general object detection & segmentation, such as PASCAL VOC (Everingham et al., 2010) and MS-COCO (Lin et al., 2014), which contain tens to hundreds of thousands of precisely annotated images, existing COD datasets (Fan et al., 2020; Lv et al., 2021) typically comprise only a few thousand samples. This scarcity arises from two intrinsic challenges: (i) camouflaged objects are inherently difficult to annotate due to their high visual similarity to backgrounds, and (ii) camouflaged scenes with high quality are themselves rare and costly to capture in the real world. As a result, data insufficiency has become a critical bottleneck for improving both the robustness and generalization of COD models.
@@
Recent advances in image generation models, including GAN (Goodfellow et al., 2020) and diffusion models (Ho et al., 2020; Rombach et al., 2022), offer a promising alternative, while leveraging high-quality synthetic data to augment training has been shown to be an effective and cost-efficient strategy in multiple vision tasks such as salient object detection (Kupyn et al., 2025). This observation motivates the camouflaged image synthesis task, whose core objective is to generate realistic and diverse camouflaged scenes that can meaningfully complement limited real-world datasets and thereby improve downstream COD performance.
@@
Figure 2: The proposed KL divergence-based metric. Existing generative evaluation metrics, such as FID and KID, are insufficient to measure the similarity between foreground and background, while our proposed $\mathrm{KL}_{BF}$ provides a quantitative assessment of camouflage. Top: Image with a higher $\mathrm{KL}_{BF}$. Bottom: Our generated image with a lower $\mathrm{KL}_{BF}$, more camouflaged.
@@
Why Out-Painting: Existing approaches to camouflaged image synthesis can be broadly categorized into two paradigms. The first paradigm, exemplified by style transfer-based methods such as LCGNet (Li et al., 2022b), modifies the appearance of a target object to match a given background. However, this process often alters object-specific textures and visual cues to such an extent that the semantic identity of the target is partially or entirely compromised. Consequently, these methods achieve camouflage by enforcing visual uniformity, rather than by modeling the natural interaction between an object and its environment.
@@
The second paradigm, represented by out-painting-based methods such as LAKE-RED (Zhao et al., 2024), follows a fundamentally different strategy. Instead of altering the target object, these methods preserve its semantic content and generate surrounding background regions whose textures and structures are consistent with the target appearance. This formulation enables camouflage to emerge from contextual compatibility between the object and its environment, rather than from direct manipulation of the object itself.
@@
Compared with style transfer-based approaches, out-painting offers several advantages. First, it maintains the semantic integrity of the target while avoiding the need for an externally provided background image. More importantly, it aligns more closely with the natural principles of camouflage observed in the real world. Camouflage is inherently a relative and environment-dependent phenomenon: the same object may appear well camouflaged in one context but highly salient in another (Wang et al., 2025). From the perspective of natural selection, creatures do not actively redesign their appearance to match each environment; instead, environmental variation favors the survival of individuals whose appearances are already compatible with their surroundings. This perspective naturally corresponds to the out-painting formulation.
@@
Why Background Control: For out-painting-based methods to produce realistic and effective camouflaged training data, we argue that two key requirements must be simultaneously satisfied. First, the generated background should exhibit strong appearance compatibility with the target, particularly in terms of color distribution and texture patterns, so that the target is genuinely camouflaged under visual perception. Second, the background must convey semantically plausible content consistent with real-world priors; for instance, aquatic objects should appear in water rather than in visually incompatible contexts.
@@
(1) Controlled camouflaged image synthesis. To overcome the limitations of existing out-painting-based methods in background generation, we introduce a novel out-painting-based framework termed RealCamo, which integrates explicit layout controls and textual-visual guidance. We argue that background layout plays a critical role in determining semantic plausibility. Building upon the contrast control in CamoAny, we incorporate additional depth and holistically-nested edge information into ControlNet to explicitly constrain the spatial layout of generated scenes, thereby preserving semantic realism. Furthermore, to enhance camouflage effectiveness at the texture level, we propose a textual–visual guidance mechanism that combines a unified fine-grained textual task description with additional visual cues retrieved according to target texture characteristics, enhancing visual fidelity and realism of generation.
(2) Camouflage effectiveness evaluation metrics. A realistic camouflaged image should preserve the original scene layout while achieving camouflage through modifications in background color and texture. Existing generative evaluation metrics, such as FID and KID, are insufficient to capture this requirement. We therefore introduce SSIM to quantify structural consistency between generated and original images as an indicator for semantic realism. In addition, we propose a KL divergence-based metric $\mathrm{KL}_{BF}$, which measures the similarity between background and foreground pixel distributions, providing a principled quantitative assessment of camouflage effectiveness, as shown in Fig. 2.
(3) Downstream evaluation. To assess the practical utility of the generated camouflaged images, we construct a large-scale synthetic dataset, SynCOD12K, as supplementary training data for camouflaged object detection (COD). Extensive experiments on four COD benchmarks demonstrate consistent performance improvements, with further investigations on the effects of camouflage-augmented data on related salient and general object detection tasks.
@@ 2 Related Work
Synthetic Dataset Generation. Synthetic data has become an effective strategy for alleviating data scarcity in computer vision (Mumuni et al., 2024). Advances in generative models, including GANs and diffusion models, have enabled the construction of large-scale, high-quality synthetic datasets. DatasetGAN (Zhang et al., 2021) generates image-annotation pairs by exploiting the latent feature space of pretrained GANs, while BigDatasetGAN (Li et al., 2022a) extends this paradigm to ImageNet-scale class diversity using VQGAN (Esser et al., 2021). More recently, diffusion-based methods such as DiffuMask (Wu et al., 2023a) and DatasetDM (Rombach et al., 2022) leverage text-to-image models to produce photorealistic images and corresponding annotations for downstream perception tasks. Despite their success, these approaches primarily focus on generic image synthesis and annotation, and remain limited when applied to structurally complex and semantically subtle scenarios such as camouflaged image generation.
@@
Camouflaged Image Generation. Camouflaged image generation, introduced by Chu et al. (2010), can be categorized into two paradigms: Style transfer-based background-guided approaches modify foreground appearance to blend targets into given backgrounds. DCI (Zhang et al., 2020) employs an attention-aware camouflage loss to suppress salient cues while preserving subtle perceptual features, and LCGNet (Li et al., 2022b) achieves extreme camouflage by fusing high-level foreground and background representations. In contrast, out-painting-based foreground-guided methods generate backgrounds conditioned on the foreground target. LAKE-RED (Zhao et al., 2024) introduces a knowledge retrieval–augmented framework using a pretrained codebook, while FACIG (Chen et al., 2025) refines feature integration to reduce foreground distortion. CamoAny (Das and Gopalakrishnan, 2025) further improves background synthesis through controlled out-painting and representation engineering. Despite these advances, existing methods still struggle to simultaneously achieve strong visual similarity and semantic consistency. More recently, the text-driven method CT-CIG (Qian et al., 2025) has shown promise in realistic camouflage synthesis, but the inability to provide aligned annotations limits its applicability to downstream camouflaged object detection applications.
@@
Controlled Image Generation. Recent advances in text-to-image (T2I) generation integrate large-scale text encoders, such as CLIP (Kim et al., 2021) and T5 (Raffel et al., 2020), into diffusion models, yielding powerful T2I models. To enable fine-grained spatial control, ControlNet (Agrawala, 2023) introduces an auxiliary trainable branch to inject conditional signals into pretrained diffusion models. Building upon this design, UniControl (Qin et al., 2023) employs a task-aware hypernetwork to modulate zero convolutions of the ControlNet in handling multiple conditions, while CtrLoRA (Xu et al., 2025) further reduces parameter overhead through condition-specific LoRA (Hu et al., 2022) fine-tuning, and Sharma et al. (2025) proposes an N-channel ControlNet to improve object preservation, prompt alignment, and aesthetic quality under multiple control settings.
@@ 3.1 Preliminary
Our proposed method is based on the classic Latent Diffusion Models (LDMs) (Rombach et al., 2022) and ControlNet (Agrawala, 2023). Here, we briefly revisit their basics, with more details can be found in Appendix B.
Figure 3: The pipeline of our camouflaged image generation framework RealCamo. Explicit layout controls (i.e., the contrast, depth, and HED controls) are extracted by the Layout Controls Generator and injected into ControlNet, while the multimodal condition produced by the Textual-Visual Condition Generator is used as guidance in refining the camouflage generation process.
Latent Diffusion Models. LDMs generate high-quality images by reversing a Markov forward process in the latent space. Specifically, Given a VAE (Rezende et al., 2014) with encoder $\mathcal{E}$ and decoder $\mathcal{D}$, Gaussian noise $\epsilon$ is progressively added to a latent representation $z_{0}=\mathcal{E}(x_{0})$ of the original image $x_{0}$, resulting in a noisy image $z_{t}=\alpha_{t}z_{0}+(1-\alpha_{t})\epsilon$, with more noise being introduced when $\alpha_{t}$ decreases as timestep $t$ increases. A denoising network $\epsilon_{\theta}$ is trained to predict the added noise given a condition $c^{t}$ following:
@@
ControlNet. ControlNet extends the structural controllability of LDMs by incorporating additional conditions ($c^{f}$) like Canny edges and human pose. With the introduction of an additional trained branch aided by zero convolutions, ControlNet reformulates Equation 1 as:
@@ 3.2 Framework Overview
As illustrated in Fig. 3, the proposed RealCamo framework is built upon Stable Diffusion 1.5 (Rombach et al., 2022) and ControlNet (Agrawala, 2023), enabling controllable camouflaged image generation through complementary structural and semantic conditioning. Control signals are injected in two ways: (i) structural constraints introduced via ControlNet to regulate spatial layout, and (ii) multimodal prompts injected through cross-attention to guide image content. These mechanisms align with our objective that high-quality camouflage requires both visual similarity and semantic consistency between foreground targets and background scenes. Accordingly, we employ a Layout Controls Generator to provide explicit structural priors via ControlNet, and a Textual–Visual Condition Generator that integrates a unified fine-grained task description with background images retrieved by a Texture-Oriented Background Retrieval module, producing informative multimodal prompts to effectively steer the camouflage generation process.
@@ 3.3 Layout Controls Generator
We argue that semantic consistency in camouflaged image generation is largely determined by the structural layout of the image. Prior methods (Zhao et al., 2024; Das and Gopalakrishnan, 2025) often yield cluttered or physically implausible backgrounds due to the lack of explicit layout constraints. To address this, we propose a Layout Controls Generator (LCG) that extracts complementary structural cues and injects them via ControlNet, as shown in Fig. 3. Specifically, for a given image $\mathrm{I}$ and a corresponding mask $\mathrm{M}$ which indicates the foreground target, we employ LCG to extract contrast, depth, and holistically-nested edge (HED) controls as explicit structural priors to promote coherent spatial organization and improved foreground–background semantic alignment:
$c_{\mathrm{cst}},c_{\mathrm{dep}},c_{\mathrm{hed}}=LCG(\mathrm{I},\mathrm{M},is\_training).$ (3)
Each control signal is embedded using a dedicated MLP (Qin et al., 2023) due to distributional differences, and then fused as structural guidance for ControlNet to process.
Contrast Control. Following CamoAny, contrast control $c_{\mathrm{cst}}$ aims to preserve the foreground target while providing basic regulation of color contrast, which varies when training and inference based on is_trainingis\_training.
Depth Control. Depth control $c_{\mathrm{dep}}$ is generated using Depth Anything V2 (Yang et al., 2024), and is intended to constrain the spatial relationships within the image.
HED Control. HED control $c_{\mathrm{hed}}$ is obtained by deploying Holistically-nested edge detection (Xie and Tu, 2015) to regulate the contours of the main elements in the image.